# The Strong Birthday Problem Revisited

### v1

@@ 2.3 Associated Stirling Numbers of the Second Kind
We can write a recurrence for $n\brace k$} as follows (see graham94 page 259):
@@ 1. Case 1:
    (Forms a partition). $n$th object’s insertion creates a partition of size $r$. The number of ways in which r1r-1 objects can be chosen from n1n-1 objects is $n-1\choose r-1$, to which the $n$th object is added. The remaining nrn-r objects can be partitioned into k1k-1 partitions of size at least $r$ in ${n-r\brace k-1}_{\geq r}$ ways. Therefore, the number of ways are ${n-1\choose r-1}{n-r\brace k-1}_{\geq r}$.
@@ 2. Case 2:
    (Does not form a partition). $n$th objects does not form a new partition. Then, there are ${n-1\brace k}_{\geq r}$ ways of partitioning n1n-1 objects into $k$ partitions of size at least $r$. $n$th objects can go into any of these $k$ partitions. Therefore, the number of ways are $k{n-1\brace k}_{\geq r}$.
@@
For $r=1$, the above recurrence reduces to Stirling numbers of the second kind. For $r=2$, the above recurrence reduces to
${n\brace k}_{\geq 2}$ $=$ $k{n-1\brace k}_{\geq 2}+(n-1){n-2\brace k-1}_{\geq 2}.$
See Comtet1974 ; Howard1980 ; Zhao2008 ; Komatsu2015 for more on the associated Stirling numbers of the second kind and related identities. We will use the above-derived recurrence for the associated Stirling numbers of the second kind to formulate a recurrence for SBP.
2.4 A Recurrence for SBP using the Associated Stirling Numbers of the Second Kind
When the partitions are labeled, as in case of the birthday problem, the number of ways of putting $n$ labeled objects into $k$ labeled partitions such that each partition has at least $r$ objects in it can be obtained by considering all $k!$ permutations of the $k$ labels designating the partitions. Therefore,
@@ 2.4 A Recurrence for SBP using the Associated Stirling Numbers of the Second Kind
$=$ $\begin{cases}0,&\text{if $(n\leq 0\land k>0)\lor(n>0\land k\leq 0)\lor(n<kr)$};\\ 1,&\text{if $n=0\land k=0$};\\ k!\left(k{n-1\brace k}_{\geq r}+{n-1\choose r-1}{n-r\brace k-1}_{\geq r}\right),&\text{otherwise.}\end{cases}$ (14)
This recurrence has a nice interpretation as one more variant of the birthday problem – how many ways are there so that $n$ people’s birthdays can be assigned to $m$ days of a year so that each day has $r\geq 2$ person’s birthdays assigned to it, and the corresponding probability problem can be stated accordingly.
@@
Alternative Recurrence for SBP: For the strong birthday problem, we can derive an alternative recurrence using the above recurrences as follows.
The number of ways in which $n$ people can have birthdays out of $m$ possible days so that each birthday must be shared by at least 2 people has the following steps in counting. The number of ways of choosing $k$ birthdays out of $m$ days is $m\choose k$}, where $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor$. Each such choice of $k$ has $T_{2}(n,k)=k!{n\brace k}_{\geq 2}$ ways of distributing $n$ people into $k$ days so that each day has at least 2 person’s birthdays incident on it. Thus, the total number of ways is given by
$N$ $=$ $\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}{m\choose k}T_{2}(n,k)=\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}{m\choose k}k!{n\brace k}_{\geq 2}.$ (18)
Since the number of ways in which $n$ people can be assigned $m$ days as birthdays is $m^{n}$, which is the size of the sample space, the probability that each person shares her/his birthday with someone in the group of $n$ people is
@@
We can have a simple generalization to the strong birthday problem when we ask the following question: What is the probability that in a group of $n$ people and $m$ days in a year, each day has at least $r\geq 2$ birthdays incident or is birthday-free? In other words, what is the probability that each person’s birthday is shared by at least r1r-1 other people for $r\geq 2$? The solution can be written straightforwardly, by observing that $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{r}\end{array}\!\!\right\rfloor$, and is given by
@@
The above formula can be computed using dynamic programming. The outer summation takes $\operatorname{O}\left(n/r\right)$ time, and there are $\operatorname{O}\left(n^{2}/r\right)$ number of unique subproblems. Computation of the binomial coefficient $m\choose k$} and $k!$ takes $\operatorname{O}\left(n/r\right)$ time with memoization. Again assuming that per state we use only constant time (by disregarding the bit complexity of large number addition and multiplication). Fixing $r=2$, the overall time complexity is $\operatorname{O}\left(n^{3}\right)$ and the space complexity can easily be shown to be $\operatorname{O}\left(n^{2}\right)$.
The probability computation using the recurrence in Equation 22 is has better time and space complexities than the recurrence in Equation 9 we derived in Subsection 2.2.
3 Implementation and Numerical Results
We have implemented the combinatorial formula (Equation 3) and the above recurrences in Equation 9 and Equation 22 to compute the minimum number of people $n$ required for different values of the number of days $m$ so that each person has a shared birthday with someone in the group with probability at least 1/2 (see row 2 of Table 1) and at least 0.999 (see row 3 of Table 1). The results matched in all three cases, thereby proving the validity of our approaches.
$m$ 10 50 100 200 364 365 366 400 500 1000
$n_{@prob\geq 0.5}$ 41 304 690 1541 3054 3064 3073 3399 4375 9528
$n_{@prob\geq 0.999}$ 112 665 1410 2975 5653 5669 5686 6253 7937 16619
@@ 3 Implementation and Numerical Results
Table 1: Minimum number of people $n$ required for a given $m$ so that everyone has a shared birthday with someone in the group with probability at least 1/2 (row 2) and at least 0.999 (row 3).
For $n>200$, since the recurrence in the Equation 9 took much longer than the recurrence in the Equation 22 due to its higher time and space complexities. So our results for $n>200$ are computed only using Equation 3 and Equation 22. We use an arbitrary precision arithmetic library so that large integer operations can be performed. Overall, while the recurrence in Equation 22 is efficient for practical computation, we believe that the recurrence that led to Equation 9 is more intuitive and can just be derived from a sound understanding of the problem statement.
@@ 4 Significance and Extensions
As discussed in DasGupta2005 , the strong birthday problem has applications in fingerprint matching in criminology and sociology. There can be more applications in areas like job scheduling, such as modeling the queues of customers in a retail shopping store, where we might want to estimate the number of customers so that each queue has size $r\geq 2$ with a certain probability threshold. Such estimation may help in opening more queues or merging queues.
In this article, we also discussed natural extensions of the strong birthday problem which are interesting in their own right. Fast approximation algorithm to compute the values of $n$ given $m$ and a probability threshold, that can scale to arbitrary large numbers is an interesting direction to pursue. Restrictions to the uniform probability assumption, and cases where each person has a different probability to be incident on a day of a year, may have interesting properties to explore.

### v2

@@ 2.4.1 Associated Stirling Numbers of the Second Kind
We can write a recurrence for $n\brace k$} as follows (see Graham et al. (1994) page 259):
@@ 1. Case 1:
    Forms a new partition. The $n$th object’s insertion creates a new partition of size exactly $r$. The number of ways in which r1r-1 objects can be chosen from n1n-1 objects is $n-1\choose r-1$, to which the $n$th object is added. The remaining nrn-r objects can be partitioned into k1k-1 partitions of size at least $r$ in ${n-r\brace k-1}_{\geq r}$ ways. Therefore, the number of ways are ${n-1\choose r-1}{n-r\brace k-1}_{\geq r}$.
@@ 2. Case 2:
    Does not form a new partition. The $n$th object joins an existing partition. Then, there are ${n-1\brace k}_{\geq r}$ ways of partitioning n1n-1 objects into $k$ partitions of size at least $r$. The $n$th objects can go into any of these $k$ partitions. Therefore, the number of ways are $k{n-1\brace k}_{\geq r}$.
@@
For $r=1$, the above recurrence reduces to Stirling numbers of the second kind (see Equation 20). For $r=2$, which is relevant to SBP, the above recurrence reduces to
${n\brace k}_{\geq 2}$ $=$ $k{n-1\brace k}_{\geq 2}+(n-1){n-2\brace k-1}_{\geq 2}.$ (23)
See Comtet (1974); Howard (1980); Zhao (2008); Komatsu et al. (2015) for more on the associated Stirling numbers of the second kind and related identities. We will use the above-derived recurrence for the associated Stirling numbers of the second kind to formulate a recurrence for SBP.
2.4.2 A Recurrence for SBP Using the Associated Stirling Numbers of the Second Kind
When the partitions are labeled, as in case of the birthday problem where days are distinguishable, the number of ways of putting $n$ labeled objects into $k$ labeled partitions such that each partition has at least $r$ objects can be obtained by considering all $k!$ permutations of the $k$ partition labels. Therefore,
@@ 2.4.2 A Recurrence for SBP Using the Associated Stirling Numbers of the Second Kind
$=$ $\begin{cases}0,&\text{if $(n\leq 0\land k>0)\lor(n>0\land k\leq 0)\lor(n<kr)$};\\ 1,&\text{if $n=0\land k=0$};\\ k!\left(k{n-1\brace k}_{\geq r}+{n-1\choose r-1}{n-r\brace k-1}_{\geq r}\right),&\text{otherwise.}\end{cases}$
This recurrence has a nice interpretation as one more variant of the birthday problem: How many ways are there so that $n$ people’s birthdays can be assigned to $k$ days so that each day has $r\geq 2$ person’s birthdays assigned to it? The corresponding probability problem can be stated accordingly.
@@
For the strong birthday problem, we can derive a recurrence using the above recurrences in Equation 25, by counting no singleton birthdays, i.e., the number of favorable outcomes, using the following two steps.
  1. Step 1:
    Choose $k$ specific days from $m$ days. The number of ways of choosing $k$ specific days from $m$ days is $\binom{m}{k}$, where $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor$.
  2. Step 2:
    Assign $n$ People to $k$ labeled days. The number of ways of assigning $n$ people’s birthdays to $k$ specific days such that each day has $\geq 2$ person’s birthdays incident on it is $T_{2}(n,k)=k!{n\brace k}_{\geq 2}$.
Therefore, the total number of ways is given by
$N$ $=$ $\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}\binom{m}{k}T_{2}(n,k)=\sum_{k=1}^{\left\lfloor\!\!\begin{array}[]{l}\frac{n}{2}\end{array}\!\!\right\rfloor}\binom{m}{k}k!{n\brace k}_{\geq 2}.$ (28)
Since the total number of ways in which $n$ people can be assigned $m$ days as birthdays is $m^{n}$, which is the size of the sample space, the probability that each person shares her/his birthday with someone in the group of $n$ people is
@@ 2. Step 2:
We can have a simple generalization to the strong birthday problem when we ask the following question: What is the probability that in a group of $n$ people and $m$ days in a year, each day has at least $r\geq 2$ birthdays incident on it or is birthday-free? In other words, what is the probability that each person’s birthday is shared by at least r1r-1 other people for $r\geq 2$? The solution can be written straightforwardly, by observing that $1\leq k\leq\left\lfloor\!\!\begin{array}[]{l}\frac{n}{r}\end{array}\!\!\right\rfloor$, and is given by
@@
The recurrence in Equation 30 can be computed using dynamic programming, and has several significant advantages over the previous recurrences we derived (see Equation 11 and Equation16). The state space described by ${n\brace k}_{\geq 2}$ only depends on $n$ and is independent of $m$. The recurrence in Equation 23 comprises of two terms instead of three as in the previous recurrences and have small integer multipliers. Most importantly, since ${n\brace k}_{\geq 2}$ computation is independent of $m$, they can be precomputed for given $n$ and can be reused across different values of $m$ which only appears in the final summation phase not during the dynamic programming recursion. Finally, since the division by $m^{n}$ and multiplication by $\binom{m}{k}$ occurs in the final summation step, and the Stirling number recursion operates on values whose growth depends only on $n$ and $k$ with predictable intermediate sizes of large integers there by avoiding manipulations of enormous integers required to implement Equation 11, and also avoids the repeated division by $m$ required by to implement Equation16.
Computational Complexity: The dynamic programming implementation using top-down memoization has $O\!\left(n^{2}\right)$ states where each state requires $O\!\left(1\right)$ operations under unit-cost arithmetic model. Therefore, both time and space complexities are $O\!\left(n^{2}\right)$. The bottom-up dynamic programming implementation using (auxiliary) rolling arrays has time complexity $O\!\left(n^{2}\right)$, and space complexity $O\!\left(n\right)$ since only two rows of $O\!\left(n\right)$ states need to be stored to compute the current row.
2.5 Summary of Complexity of Algorithms for SBP
Table 1 summarizes the asymptotic time and space complexities of the various methods derived for solving SBP. Compared to the implementations of the recurrences in Equation 15 and Equation 18, which require $O\!\left(n^{3}\right)$ time, the recurrence in Equation 30 is asymptotically faster with a time complexity of $O\!\left(n^{2}\right)$. This efficiency gain extends to memory usage as well. The top-down memoization of the former recurrences requires $O\!\left(n^{3}\right)$ space, which can be reduced to $O\!\left(n^{2}\right)$ using bottom-up dynamic programming with rolling arrays. In contrast, Equation 30 requires only $O\!\left(n^{2}\right)$ space with top-down memoization, which further reduces to $O\!\left(n\right)$ space when implemented via bottom-up dynamic programming using rolling arrays.
Table 1: Complexity Comparison of Algorithms for SBP
Algorithm Time Complexity Space Complexity
Combinatorial Formula (Equation 9) $O\!\left(n\right)$ $O\!\left(1\right)$
Counting Recurrence (Equation 15)
   Memoization $O\!\left(n^{3}\right)$ $O\!\left(n^{3}\right)$
   Bottom-Up DP $O\!\left(n^{3}\right)$ $O\!\left(n^{2}\right)$
Probability Recurrence (Equation 18)
   Memoization (float) $O\!\left(n^{3}\right)$ $O\!\left(n^{3}\right)$
   Bottom-Up (float) $O\!\left(n^{3}\right)$ $O\!\left(n^{2}\right)$
Stirling Recurrence (Equation 30)
   Memoization $O\!\left(n^{2}\right)$ $O\!\left(n^{2}\right)$
   Bottom-Up DP $O\!\left(n^{2}\right)$ $O\!\left(n\right)$
3 Results and Discussion
The different algorithms for computing SBP were implemented in Python 3.x. To facilitate rigorous verification, the combinatorial formula, the counting-based recurrence, and the associated Stirling number recurrence were implemented using the mpmath library with 1000 decimal places of precision (mp.dps = 1000), producing identical results across all test cases. In contrast, our implementation of the probability-based recurrence utilizes standard 64-bit IEEE 754 floating-point arithmetic. This approach demonstrates a significant performance gains and exhibits lower sensitivity to rounding errors compared to counting-based methods, making it ideal for scenarios where standard precision suffices. All experiments were conducted on a macOS-based system.
Table 2 summarizes the asymptotic performances and numerical characteristics of the combinatorial formula and the optimized bottom-up dynamic programming algorithms. While the combinatorial formula offers the lowest theoretical time complexity, its reliance on arbitrary-precision libraries for stability makes the probability-based recurrence more practical for simulations using native floating-point hardware. Stirling recurrence (Equation 30) offers best of both the worlds with arbitrary precision. Although it uses mpmath arithmetic functions, due to its cleaner state space than the other two methods, and better asymptotic time and space complexities compared to the counting and probability recurrences, it provides the best balance when high precision is needed. Furthermore, Stirling recurrence reveals connections to partition theory through associated Sterling numbers of the second kind which is exploited to derive the recurrence.
Table 2: Comparison of Algorithms for SBP: Complexity, Precision, and Stability
[…]
Combinatorial Formula (Equation 9) High (mpmath) $O\!\left(n\right)$ $O\!\left(1\right)$ Low*
Counting Recurrence (Equation 15) High (mpmath) $O\!\left(n^{3}\right)$ $O\!\left(n^{2}\right)$ Moderate
Probability Recurrence (Equation 18) Standard $O\!\left(n^{3}\right)$ $O\!\left(n^{2}\right)$ High
Stirling Recurrence (Equation 30) High (mpmath) $O\!\left(n^{2}\right)$ $O\!\left(n\right)$ Moderate
$\dagger$ Space complexity reflects optimized bottom-up dynamic programming implementations.
* Requires arbitrary precision (mpmath) due to large intermediate values like $m^{n}$ and $n!$.
We selected 6 test cases to examine two critical probability thresholds across different scales of $m$:
  • Cases approaching $\approx 50\%$ probability: $(m,n)=(10,41),(50,304),(100,690)$; and
  • Cases approaching $\approx 99.9\%$ probability: $(m,n)=(10,112),(50,665),(100,1410)$.
Table 3 provides a comprehensive runtime comparison of the four implemented methods across varying scales of $m$ and $n$. The combinatorial formula based implementation with time complexity of $O\!\left(n\right)$ is much faster than the dynamic programming methods as expected, although numerical stability can be an issue for large values of $m$ and $n$. These results highlight the dramatic performance gains of closed-form solutions. By bypassing the computational depth of recurrences, these solutions offer both faster execution and simpler implementation paths.
Table 3: Performance Comparison of Algorithms for SBP: Runtime (Seconds)
Parameters Runtime (s)
Test $m$ $n$ Combinatorial Formula Counting Rec. Probability Rec. Stirling Rec.
(Equation 9) (Equation 15) (Equation 18) (Equation 30)
(mpmath) (mpmath) (float) (mpmath)
1 10 41 $0.004$ $0.019$ $0.002$ $0.004$
2 50 304 $0.027$ $4.047$ $0.350$ $0.165$
3 100 690 $0.066$ $42.073$ $3.242$ $0.885$
4 10 112 $0.008$ $0.061$ $0.007$ $0.022$
5 50 665 $0.061$ $9.987$ $1.034$ $0.807$
6 100 1410 $0.176$ $91.624$ $8.583$ $4.040$
Complexity $O\!\left(n\right)$ $O\!\left(n^{3}\right)$ $O\!\left(n^{3}\right)$ $O\!\left(n^{2}\right)$
Performance Ratio* $1\times$ $520.6\times$ $48.8\times$ $23.0\times$
* Performance Ratio is calculated relative to the Combinatorial Formula for Test 6.
All three recurrence relations, while computationally slower than the combinatorial formula, remain valuable for theoretical understanding. For example, the optimal substructures and the overlapping subproblems are not observed through the combinatorial formula; therefore, a limited insight into the structure of SBP is obtained from the combinatorial formula. In contrast, the three recurrences derived here show the different combinatorial recursive structures of the same problem, respectively, through counting, probability flow and establishing a profound connection to partition theory. Therefore, we use the implementation of the combinatorial formula primarily as a baseline for correctness verification.
For the test case ($m=100$ and $n=1410$), we give the speed up ratios with respect to the run-time of the combinatorial formula. Among the recurrence-based methods, the Stirling recurrence shows the most significant advantage; for $n=1410$ (Test 6), its $O\!\left(n^{2}\right)$ complexity reduces runtime to $4.04$ seconds, yielding a $22.7$-fold speedup over the $O\!\left(n^{3}\right)$ counting recurrence while maintaining high precision via mpmath. Additionally, the probability recurrence remains competitive for smaller $n$ by leveraging 64-bit floating-point arithmetic, demonstrating the trade-off between computational throughput and numerical precision.
Table 4 presents the minimum population size $n$ required to ensure that every individual shares a birthday with at least one other person with probabilities $P\geq 0.5$ and $P\geq 0.999$. Our various implementations yielded identical results across all test cases, verifying the correctness of the derived approaches. For $m>200$, the high time and space complexities of the counting and probability recurrences became computationally prohibitive; consequently, results for these larger scales were computed and cross-verified exclusively using the combinatorial formula (Equation 9) and the Stirling recurrence (Equation 30).
Notably, while the classical birthday problem requires only $n\approx\sqrt{m}$ individuals to find a single pair, the Strong Birthday Problem requires a significantly larger population, scaling approximately as $n\sim m\log m$, i.e., $n=O\!\left(m\log m\right)$, to ensure that no individual remains a singleton. This asymptotic behavior is similar to that of the Coupon Collector’s Problem Erdős and Rényi (1961); Motwani and Raghavan (1995) since the occupancy distribution Feller (1968) in both the problems uses Stirling number of the second kind (and their restrictions), which forms the basis for our recurrences.
@@ 3 Results and Discussion
Table 4: Minimum number of people $n$ required for a given $m$ such that every individual shares a birthday with at least one other person in the group with probability $P\geq 0.5$ and $P\geq 0.999$
Days ($m$) $n$ at $P\geq 0.5$ $n$ at $P\geq 0.999$
10 41 112
50 304 665
100 690 1410
200 1541 2975
364 3054 5653
365 3064 5669
366 3073 5686
400 3399 6253
500 4375 7937
1000 9528 16619
@@ 4 Significance and Extensions
Beyond its theoretical appeal, SBP has high-impact applications in diverse fields. As discussed in DasGupta (2005), the problem provides a framework for fingerprint matching in criminology and sociological modeling. In computer science, SBP is related to $k$-anonymity and honey encryption, where the elimination of singletons is required to ensure that every record in a database is indistinguishable from a larger anonymity set. In distributed systems and load balancing, SBP models the point at which no server node is left with a lone task, ensuring redundancy and mirroring across all occupied resources. Furthermore, the problem finds relevance in job scheduling and retail operations; for example, it can model customer queues in a shopping store where management seeks to estimate the population $n$ required such that each active queue has a size $r\geq 2$. Such estimations are vital for making real-time decisions regarding merging or opening new service channels.
Several natural extensions of SBP remain fertile ground for future research. A primary direction is the development of fast approximation algorithms to compute $n$ for arbitrary $m$ and probability thresholds, particularly for scales where $O\!\left(n^{2}\right)$ recurrences become computationally expensive. Additionally, relaxing the uniform probability assumption to account for skews or cases where individuals have non-identical probability distributions across the year would provide a more robust model for real-world phenomena.
Acknowledgments
I am thankful to Dr. Chittaranjan Tripathy for introducing me to discrete mathematics, birthday problem and then suggesting me that SBP can be an interesting problem to work on. I am thankful to Dr. Shashidhara K. Ganjugunte and Dr. Pouyan Shirzadian for numerous valuable comments on the manuscript.